OpenStudy (anonymous):

Solve this equation. 1/4 - x =x/8

OpenStudy (anonymous):

Add x on both sides: 1/4 = x + x/8. Make the RHS with a common denominator: 1/4 = (9/8)x. Multiply both sides by 8/9, we are left with: x = 2/9.

OpenStudy (anonymous):

\[x=2/9\]